OpenStudy (anonymous):

verify: 1- sin^2theta/1-costheta = -costheta

OpenStudy (anonymous):

sin^2theta+cos^2theta=1 1-sin^2theta=cos^2theta so cos^2theta/1-costheta=cos^2theta-costheta

OpenStudy (anonymous):

=costheta(costheta-1)

OpenStudy (anonymous):

so would your first step be to find the common denominator?

OpenStudy (a_clan):

1 - sin^2theta ----------- 1- costheta = 1 - (1- cos^2theta) ------------- (1-cos theta) = 1 - (1+cos theta)(1-cos theta) ----------------------- ( 1-cos theta) = 1 - ( 1 + COS THETA)

OpenStudy (anonymous):

my answer should be negative though, right?

OpenStudy (a_clan):

Last step is, = 1 - 1 -cos theta = -cos theta

OpenStudy (anonymous):

but don't you cross out both the 1-costheta 's so you'd be left with 1+costheta

OpenStudy (anonymous):

A clan is correct I though the problem was \[ \frac{ 1-\sin ^{2}\theta }{ 1-\cos}\]

OpenStudy (anonymous):

oh... i don't understand the last step though

OpenStudy (mathstudent55):

\( \dfrac{1 - \sin^2 \theta}{1 - \cos \theta} = - \cos \theta \) \( 1 - \sin^2 \theta = (1 - \cos \theta)(-\cos \theta) \) \( 1 - \sin^2 \theta = -\cos \theta + \cos^2 \theta \) \( \cos^2 \theta = -\cos \theta + \cos^2 \theta \) \(0 = -\cos \theta\) This is not an identity.

OpenStudy (anonymous):

the 1 is separate

OpenStudy (anonymous):

i should've made that more clear

OpenStudy (mathstudent55):

You mean 1 is not in the fraction in the numerator, but it is in the denominator.

OpenStudy (anonymous):

its neither

OpenStudy (anonymous):

1/1 - sin^2theta/1-costheta = -costheta

OpenStudy (anonymous):

^ hopefully that makes more sense

OpenStudy (mathstudent55):

I still don't underestand it. You can use the equation editor or the draw tool to make it clear.